# An RSK correspondence for cylindric tableaux

### 10 Mar 2026 v1

@@ 10.1. Wilf-equivalences
First, by combining the oscillating RS correspondence with conjugation, one can prove the shape-Wilf-equivalence of the pattern sets $\{d\cdots 1,1\cdots L\}$ and $\{L\cdots 1,1\cdots d\}$. This result is essentially stated in [20, Thm. 3], but not in the language of pattern-avoidance.
Next, it is a general fact that shape-Wilf-equivalence is well-behaved with respect to “exchanging prefixes”. In particular, the shape-Wilf-equivalence given in the previous paragraph automatically implies the shape-Wilf-equivalence of the pattern sets $\{d\cdots 1(d+1),1\cdots L(L+1)\}$ and $\{L\cdots 1(L+1),1\cdots d(d+1)\}$. This prefix-exchange property was first proved in [1, Thm. 2.3] for singleton sets of patterns, but the proof easily carries over to sets with multiple patterns (see [6, Thm. 2]). It can also be adapted to the setting of symmetric fillings of Young diagrams (see [16, Thm. 3.1]), which proves the involution case of the corollary. ∎
@@ 10.2. Asymptotics
This asymptotic also implies the much weaker estimate $|S_{n}^{(\infty,L)}|=L^{2n+o(1)}$ as $n\to\infty$. Thus, the number of permutations in $S_{n}$ that avoid $1\cdots(L+1)$ grows exponentially as $n\to\infty$, and the growth rate is $L^{2}$. In the language of pattern-avoiding permutations this growth rate is called the Stanley–Wilf limit (see [17, Sec. 6.1.4]) of the pattern $1\cdots(L+1)$.
@@
Combing this theorem with the cylindric RS correspondence, we see that $|S_{n}^{(d,L)}|$ is equal to the expression in (7). The latter expression can be viewed as the expected value of $|\mathrm{tr}(U)|^{2n}$ where $U$ is a random $d\times d$ matrix drawn from a certain discrete analogue of the circular unitary ensemble. This is exactly analogous to the theorem of Rains for $|S_{n}^{(\infty,L)}|$, and we will use this to derive the asymptotics for $|S_{n}^{(d,L)}|$ as $n\to\infty$.
@@
From Theorem 10.5 and the cylindric RS correspondence, we see that (7) is equal to $|S_{n}^{(d,L)}|$. Hence, our goal is to understand the asymptotics of (7) as $n\to\infty$. If $d$ and $L$ are fixed, then (7) can be rewritten as a sum over a fixed collection of terms of the form $Br^{2n}$, where each $B$ and $r$ depend on the values $t_{1},\ldots,t_{d}$. In particular, for a given choice of $t_{1},\ldots,t_{d}$ the values of $B$ and $r$ are given by
@@
The value of $B$ is nonzero if and only if the $t_{j}$’s are distinct. On the other hand, the value of $r$ is largest when the $e(\tfrac{t_{l}}{M})$’s are clustered together as closely as possible on the unit circle. Hence, we claim that the (constrained) maximum value of $r$ occurs precisely when the $t_{j}$’s are consecutive integers modulo $M$ rearranged in any order. To prove this one can check that for any fixed $\theta\in\mathbb{R}$ it is possible to increase the component of $\sum_{l=1}^{d}e(\tfrac{t_{l}}{M})$ in the direction of $e(\theta)$ by replacing one of the roots of unity in the sum with another root of unity that is closer to $e(\theta)$. The claim follows easily from this observation.
@@
A consequence of this Theorem 10.4 is that the Stanley–Wilf limit of the pattern set $\{d\cdots 1(d+1),1\cdots(L+1)\}$ is $\left(\frac{\sin(\pi d/M)}{\sin(\pi/M)}\right)^{2}$. Note that this quantity is unchanged when $d$ and $L$ are interchanged, which is consistent with the Wilf-equivalence established in Corollary 10.2. One can also check that $C_{d,L}=C_{L,d}$, although this is less obvious from the stated formula.
@@ 11. RS, $d$-RS, and skew $d$-RS growth diagrams
In this section we discuss an important special case of the growth diagrams considered in this paper. Namely, we consider the case where the size of adjacent partitions in these growth diagram can differ by at most 1. Any RSK, $d$-RSK, or skew $d$-RSK growth diagram that satisfies this restriction will be referred to as an RS, $d$-RS, or skew $d$-RS growth diagram, respectively.
@@
(c) The skew $d$-RS local rule. In these configurations, $\kappa$ is a $d$-staircase and $i,j\leq d$. The Elizalde local rule is the same except that $\kappa$ is a $(d,L)$-staircase and the first configuration is removed.
@@
The oscillating RS correspondence is well known (see [20, Thm. 1], [31, Ch. 4]). The oscillating skew $d$-RS correspondence was essentially constructed by Elizalde [10] as we alluded to in Section 7.4. Elizalde was interested in the special case of $(d,L)$-cylindric standard oscillating tableaux, but aside from this minor difference the configurations shown in Figure 9(c) are equivalent to those given in [10, Sec. 5].
@@
The domain of the Bloom–Saracino map is the set of all fillings whose row and column sums are at most 1. It is not bijective, but it becomes a bijection (the oscillating $d$-RS correspondence) upon restricting to fillings that avoid $d\cdots 1(d+1)$. This is not stated explicitly in [4], but it is an immediate consequence of the main result. In fact, Bloom and Saracino’s work goes further by giving a characterization of the fibers of the Bloom–Saracino map. In particular they show that the map is invariant under a certain transformation on fillings.
@@ 12. Continuous piecewise-linear analogues
The RSK-type correspondences that we have described in this paper also have continuous piecewise-linear analogues. For the classical RSK correspondence, such an analogue is well-known (see [3, Sec. 1] for a discussion of this). We shall briefly describe how these analogues work and how they carry over to $d$-RSK (or skew $d$-RSK).
@@
The definitions of the RSK, $d$-RSK, and skew $d$-RSK local rules were written in terms of continuous piecewise-linear functions. Consequently, the proofs of our main results generalize with hardly any modification. This leads to a continuous piecewise-linear analogues of the cylindric correspondences as well.

### 13 Mar 2026 v2

@@ 10.1. Wilf-equivalences
First, by combining the oscillating RS correspondence with conjugation, one can prove the shape-Wilf-equivalence of the pattern sets $\{d\cdots 1,1\cdots L\}$ and $\{L\cdots 1,1\cdots d\}$. This result is essentially stated in [21, Thm. 3], but not in the language of pattern-avoidance.
Next, it is a general fact that shape-Wilf-equivalence is well-behaved with respect to “exchanging prefixes”. In particular, the shape-Wilf-equivalence given in the previous paragraph automatically implies the shape-Wilf-equivalence of the pattern sets $\{d\cdots 1(d+1),1\cdots L(L+1)\}$ and $\{L\cdots 1(L+1),1\cdots d(d+1)\}$. This prefix-exchange property was first proved in [1, Thm. 2.3] for singleton sets of patterns, but the proof easily carries over to sets with multiple patterns (see [6, Thm. 2]). It can also be adapted to the setting of symmetric fillings of Young diagrams (see [17, Thm. 3.1]), which proves the involution case of the corollary. ∎
@@ 10.2. Asymptotics
This asymptotic also implies the much weaker estimate $|S_{n}^{(\infty,L)}|=L^{2n+o(1)}$ as $n\to\infty$. Thus, the number of permutations in $S_{n}$ that avoid $1\cdots(L+1)$ grows exponentially as $n\to\infty$, and the growth rate is $L^{2}$. In the language of pattern-avoiding permutations this growth rate is called the Stanley–Wilf limit (see [18, Sec. 6.1.4]) of the pattern $1\cdots(L+1)$.
@@
Combining this theorem with the cylindric RS correspondence, we see that $|S_{n}^{(d,L)}|$ is equal to the expression in (7). The latter expression can be viewed as the expected value of $|\mathrm{tr}(U)|^{2n}$ where $U$ is a random $d\times d$ matrix drawn from a certain discrete analogue of the circular unitary ensemble. Hence, this is an exact analogue of the theorem of Rains stated above involving $|S_{n}^{(\infty,L)}|$.
@@
As noted above, Theorem 10.5 and the cylindric RS correspondence imply that (7) is equal to $|S_{n}^{(d,L)}|$. Hence, our goal is to understand the behavior of (7) as $n\to\infty$. If $d$ and $L$ are fixed, then (7) can be rewritten as a sum over a fixed collection of terms of the form $Br^{2n}$, where each $B$ and $r$ depend on the values $t_{1},\ldots,t_{d}$. In particular, for a given choice of $t_{1},\ldots,t_{d}$ the values of $B$ and $r$ are given by
@@
The value of $B$ is nonzero if and only if the $t_{j}$’s are distinct. On the other hand, the value of $r$ is largest when the $e(\tfrac{t_{l}}{M})$’s are clustered together as closely as possible on the unit circle. Hence, it is intuitively clear that the (constrained) maximum value of $r$ should occur precisely when the $t_{j}$’s are consecutive integers modulo $M$ rearranged in any order. To give a rigorous proof of this claim, one can check that for any fixed $\theta\in\mathbb{R}$ it is possible to increase the component of $\sum_{l=1}^{d}e(\tfrac{t_{l}}{M})$ in the direction of $e(\theta)$ by replacing one of the roots of unity in the sum with another root of unity that is closer to $e(\theta)$. The claim follows easily from this observation.
@@
A consequence of this theorem is that the Stanley–Wilf limit of the pattern set $\{d\cdots 1(d+1),1\cdots(L+1)\}$ is $\left(\frac{\sin(\pi d/M)}{\sin(\pi/M)}\right)^{2}$. Note that this quantity is unchanged when $d$ and $L$ are interchanged, which is consistent with the Wilf-equivalence established in Corollary 10.2. One can also check that $C_{d,L}=C_{L,d}$, although this is less obvious from the stated formula.
@@ 11. RS, $d$-RS, and skew $d$-RS growth diagrams
In this section we discuss an important special case of the growth diagrams considered in this paper. Namely, we consider the case where the sizes of adjacent partitions in these growth diagram can differ by at most 1. Any RSK, $d$-RSK, or skew $d$-RSK growth diagram that satisfies this restriction will be referred to as an RS, $d$-RS, or skew $d$-RS growth diagram, respectively.
@@
(c) The skew $d$-RS local rule. In these configurations, $\kappa$ is a $d$-staircase and $i,j\leq d$. The Elizalde local rule is the same except that $\kappa$ is a $(d,L)$-staircase, and the first three configurations are removed.
@@
The oscillating RS correspondence is well known (see [21, Thm. 1], [32, Ch. 4]). The oscillating skew $d$-RS correspondence was essentially constructed by Elizalde [10] as we discussed in Section 7.4. Elizalde was interested in the case of tableaux of that are $(d,L)$-cylindric and standard, but aside from this minor difference the configurations shown in Figure 9(c) are equivalent to those given in [10, Sec. 5].
@@
The domain of the Bloom–Saracino map is the set of all fillings of Young diagrams whose row and column sums are at most 1. It is not a bijective map, but it becomes a bijection (the oscillating $d$-RS correspondence) upon restricting to fillings that avoid $d\cdots 1(d+1)$. This is not stated explicitly in [4], but it is an immediate consequence of the main result given there. In fact, Bloom and Saracino’s work goes further by giving a characterization of the fibers of the Bloom–Saracino map. In particular they show that the map is invariant under a certain transformation on fillings.
@@ 12. Continuous piecewise-linear analogues
The RSK-type correspondences that we have described in this paper also have continuous piecewise-linear analogues. For the classical RSK correspondence, such an analogue is well known (see [3, Sec. 1] for a discussion of this). We shall briefly describe how these analogues work and how they carry over to $d$-RSK (or skew $d$-RSK).
@@
The definitions of the RSK, $d$-RSK, and skew $d$-RSK local rules were written in terms of continuous piecewise-linear functions. Consequently, the proofs of our main results generalize with hardly any modification. This leads to continuous piecewise-linear analogues of the cylindric correspondences as well.